OpenStudy (anonymous):

can anyone help me more with equations of ellipses and hyperbolas? i would love to get passed this stuff already. been on it for a few weeks.

OpenStudy (anonymous):

I'd love to your problems of ellipses and hyperbolas.

OpenStudy (anonymous):

the problem is \[4x ^{2}+9y ^{2}-8x-18y-23=0\] find the center and separate the values with a comma

OpenStudy (anonymous):

Are you telling me that center (1,1) Major axis vertices (4,1) , (-2,1) Minor axis vertices (1,3) , (1,-1) Foci (1 + sqrt 13,1) , (1 - sqrt 13,1) were wrong??? I would like to know because I am sure it is correct.

OpenStudy (anonymous):

well on the assignment i was doing, i can only miss 2 questions. 3 were wrong. so i exited out of the assignment because i was going to fail it anyways. so then i had to start over. and now it's not always the same questions

OpenStudy (anonymous):

Well this one was the same one... I think on the second one you didn't type it in correctly. I think you missed a sign. Now when it comes to the sqrt 13, I don't know if it want you to use a calculator and estimate or use the sqrt 13.. You will have to know which one they want.

OpenStudy (anonymous):

I will stick with you until you get enough correct if you want. If not, I will move on

OpenStudy (anonymous):

you have to figure out the square & round it. but yeah, it's the same one. but it's very uncommon that it will repeat questions. i have no idea at all how to do any of this. kinda the reason why i have been on it for almost a month now

OpenStudy (anonymous):

Put those answers in and tell me if they were wrong.

OpenStudy (anonymous):

i just brought the assignment back up, it's not the same question.

OpenStudy (anonymous):

What is the question?

OpenStudy (anonymous):

find the center: \[-4x ^{2}+y ^{2}-8x-12y+36=0\] use commas to separate the values

OpenStudy (anonymous):

How about (-1,6) .. does it say it is an hyperbola?

OpenStudy (anonymous):

yeah, it's a hyperbola

OpenStudy (anonymous):

Next question. find the vertices for the hyperbola. write your answer as (x1,y1),(x2,y2)

OpenStudy (anonymous):

(-2,6),(0,6)

OpenStudy (anonymous):

find the foci for this hyperbola write your answer as (x1,y1),(x2,y2) round if necessary to the nearest hundredth.

OpenStudy (anonymous):

(-3.24,6),(1.24,6)

OpenStudy (anonymous):

find the equation that represents the asymptotes of this hyperbolas A.\[y=\pm2x\] B.\[y-6=\pm1/2(x+1)\] C.\[y-6=\pm2(x+1)\] D.\[y=\pm1/2x\]

OpenStudy (anonymous):

A.

OpenStudy (anonymous):

WAIT

OpenStudy (anonymous):

Did you put it in... I don't think it is correct. I forgot about the center..

OpenStudy (anonymous):

no not yet

OpenStudy (anonymous):

It is C.

OpenStudy (anonymous):

find the center of this ellipse: \[9x ^{2}+y ^{2}+36x-4y+31=0\] use commas to separate the value

OpenStudy (anonymous):

(-2,2)

OpenStudy (anonymous):

find the major axis vertices for the ellipse write your answer in this form (x1,y1),(x2,y2)

OpenStudy (anonymous):

(-2,5),(-2,-1)

OpenStudy (anonymous):

find the major axis vertices for this ellipse write your answer in this form (x1,y1),(x2,y2)

OpenStudy (anonymous):

minor not major

OpenStudy (anonymous):

(-3,2),(-1,2)

OpenStudy (anonymous):

is that the minor?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

find the foci vertices for this ellipse write your answer in this form (x1,y1),(x2,y2) round if necessary to the nearest hundredth

OpenStudy (anonymous):

(-2,5.16),(-2,-1.16)

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

U R WELCOME